-2[-2-3(-2-2)][-2(-4)-3]
Answer by math-vortex(648)

Simplify -2[-2-3(-2-2)][-2(-4)-3].
------------------------------------
Hi, there--
.
You will need to use order of operations to simplify this expression. Let's begin simplifying inside the parentheses.
.
-2[-2-3(-2-2)][-2(-4)-3]=
.
(-2-2)=-4, so our expression becomes
-2[-2-3(-4)][-2(-4)-3]=
.
Now we can clear the parentheses. Remember the -3(-4) is a multiplication, -3(-4)=12. Likewise, -2(-4)=8. Are expression has been simplified to,
-2[-2+12][8-3]=
.
Now we simplify the expressions inside the brackets.
-2[10][5]=
.
Finally, we multiply the remaining three numbers together.
-2[50]=
-100
.
I hope this helps! Feel free to email, if any part of this explanation is unclear.
.
Ms.Figgy
math.in.the.vortex@gmail.com